# Biblio.bib

### v2

@@ 2.1.1. Tensorial analysis
(2.1) $\mathrm{Sym}:V^{\otimes m}\to S^{m}V\quad e_{i_{1}}\otimes e_{i_{2}}\otimes\cdots\otimes e_{i_{m}}\mapsto\frac{1}{m!}\sum_{\sigma\in\mathfrak{S}_{m}}e_{i_{\sigma(1)}}\otimes e_{i_{\sigma(2)}}\otimes\cdots\otimes e_{i_{\sigma(m)}}$
@@ 2.1.2. Analysis on $SM$
We consider the unit tangent bundle $SM:=\{(x,v)\in TM\mid g(v,v)=1\}$. We will write $\varphi_{t}$ for the geodesic flow. The geodesic vector field is then the generator of the flow, i.e $X=\tfrac{d}{dt}\varphi_{t}|_{t=0}.$ We denote by $\pi:SM\to M$ the canonical projection on the base. The vertical distribution is defined to be
@@
where $\mathcal{P}_{t}$ denotes the parallel transport along the geodesic. We recall that the metric $g$ on the base induces naturally a metric $g_{\text{Sas}}$ (called the Sasaki metric, see [Pat, Chapter 1]) on the unit tangent bundle $SM$. We have the following orthogonal splitting:
@@
[…]
@@
Consider a local frame $(e_{1},e_{2},\ldots,e_{n})$ parallel at $x_{0}$ (i.e $\nabla_{e_{i}}e_{j}=0$ at $x_{0}$). Recall that the action of $\nabla^{*}$ on a $m$-tensor $T$ can be described locally by:
@@
where $(R(e_{i},v)v)^{\mathbb{V}}\in\mathbb{V}$ is the vertical component of $R(e_{i},v)v$ (in other words, one has $\mathcal{K}(x,v)(R(e_{i},v)v)^{\mathbb{V}}=R(e_{i},v)v$) and the computation of $[X,X_{j}]$ can be found in the proof of [Lef, Lemma 13.1.6]. This means that the expression above simplifies into
@@
We will write $X=X_{+}+X_{-}$, where $X_{\pm}:C^{\infty}(M,\Omega_{k})\to C^{\infty}(M,\Omega_{k\pm 1})$, see [Lef, Lemma 14.2.1].
@@
We have the important relation:
@@ 2.1.4. Pestov’s identity
We introduce an ”energy estimate” due to Pestov (and first written in coordinate-free way in [Kni, Appendix Theorem 1.1], see also [PSU, Corollary 3.6]). It will be crucial in the proof of solenoidal injectivity of Proposition 4.1. More precisely, we will need the following estimate which is implied in negative curvature by the localized Pestov identity. We will use the result of [Lef, Theorem 14.3.4], which is valid in dimension $n\geq 4$:
@@
A consequence is the non-existence of conformal Killing fields of degree $m\geq 1$ ([Lef, Lemma 16.1.5]):
@@
Finally, let $S\in C^{\infty}(M;S^{2}T^{*}M)$ be a divergence-free tensor, i.e $D_{g_{0}}^{*}S=0$ and write $S=S_{0}+L(h)$ with $S_{0}\in C^{\infty}(M;S_{0}^{2}T^{*}M)$ and $h\in C^{\infty}(M)$. Then
$D_{g_{0}}^{*}S=0\iff D_{g_{0}}^{*}(S_{0}+L(h))=0\iff D_{g_{0}}^{*}S_{0}+D_{g_{0}}h=0,$
where we have used the relation $D_{g_{0}}^{*}L(h)=D_{g_{0}}h$ on 00-tensors. The divergence operator corresponds to a rescaled version of $X_{-}$. Indeed, using (2.10) and (2.7), one has, for any $S_{0}\in C^{\infty}(M;S^{2}_{0}T^{*}M)$ and $p\in C^{\infty}(M;T^{*}M)$,
@@
In particular, the functional $\Phi$ is constant along any orbit $\mathcal{O}(g)$ and the kernel of the Hessian $d^{2}\Phi_{g_{0}}$ always contains the tangent space $T_{g_{0}}\mathcal{O}(g_{0})$. We will hence need to prove an injectivity result of the Hessian on a ”transverse slice” to $T_{g_{0}}\mathcal{O}(g_{0})$. We remark the following fact:
@@
in particular, a natural transverse slice is provided by the kernel of the adjoint $D_{g_{0}}^{*}$, [Lef, Theorem 14.1.8].
@@
The analysis will take place on the space of divergence-free (or solenoidal) tensors $C^{\infty}(M;S^{m}T^{*}M)\cap\mathrm{Ker}(D_{g_{0}}^{*})$. We thus recall the following lemma which allows to ”project” a metric $g$ on solenoidal tensors. It was obtained in this form in [GuKnLef, Lemma 2.4] but the idea goes back to Ebin [Eb].
@@
Let $k\geq 2$, and $\alpha\in(0,1)$. Then there exists a neighborhood $\mathcal{U}$ of $g_{0}$ in the $C^{k,\alpha}$-topology such that for any $g\in\mathcal{U}$, there is a unique $\phi\in\mathrm{Diff}_{0}(M)$ of regularity $C^{k+1,\alpha}$, close to identity, such that $\phi^{*}g\in\mathrm{Ker}(D_{g_{0}}^{*})$ is divergence free. Moreover, there exists $\varepsilon>0$ and $C>0$ such that
$\|g-g_{0}\|_{C^{k,\alpha}}\leq\varepsilon\ \Rightarrow\ \|\phi^{*}g-g_{0}\|_{C^{k,\alpha}}\leq C\|g-g_{0}\|_{C^{k,\alpha}}.$
@@ 2.2. The operator $\Pi$
[…]
Then for any smooth function $f\in C^{\infty}(SM)$ with zero mean, we define
[…]
The second equality follows from the exponential decay of correlations [Li], see for instance [GuKnLef]. The Hessian $d^{2}\Phi(g_{0})$ will be rewritten using the operator $\Pi$ in Theorem 6.
A more microlocal definition of the operator $\Pi$ was given originally in [Gu] which allowed to track its analytical properties but we will not need to recall this definition here. We now define
@@
The operator $\Pi_{m}$ was shown to be a pseudo-differential operator of order $-1$ which is elliptic and injective on solenoidal tensors (see [Gu, GuLef, Theorem 3.5, Lemma 4.3]). In negative sectional curvature, we have the following coercive estimate, which we will use crucially to obtain the stability estimate (1.3).
@@ 2.3. Locally symmetric spaces
Let $M^{n}$ be a differentiable manifold. A metric $g$ on $M$ is locally symmetric if its curvature tensor $R_{g}$ is symmetric : $\nabla_{g}R_{g}=0$.
We will use a more algebraic characterization of locally symmetric spaces. Indeed, it is known that [Most, Chapter 19] compact locally symmetric spaces of negative sectional curvature are obtained as compact quotients of hyperbolic spaces. More precisely, one has
Proposition 2.5.
@@
We note that for a locally symmetric metric, the Lyapunov exponents of the geodesic flow are independent of the base point and given by:
@@
More precisely, if $(x,v)\in SM$ and $Y$ is the corresponding geodesic vector field, then one can find parallel orthonormal fields $Y_{1}=Y,Y_{2},\ldots,Y_{n}$ (along the geodesic defined by $(x,v)$) such that $R(Y,Y_{j})Y=-\lambda_{j}^{2}Y_{j}$ with $\lambda_{1}=0$, $\lambda_{2}=\ldots=\lambda_{r}=2$ and $\lambda_{r+1}=\ldots=\lambda_{n}=1$ (for $r=1,2,4,8$ depending on the type of locally symmetric space). We have
@@
where $Z\mapsto R(Y_{j},Z)Y_{j}$ is a symmetric endomorphism. Moreover, since one has $-4\leq g(R(Y_{j},Z)Y_{j},Z)\leq-1$ for any unit length $Z$, this implies that
@@
This means that the unstable Jacobian $J^{u}(v):=-\tfrac{d}{dt}|_{t=0}\mathrm{det}(d\varphi_{t}(v)|_{E_{u}(v)})$, where $E_{u}(v)$ is the unstable bundle at $v\in SM$, is a constant function of $v$. In particular, its equilibrium state (the Sinai-Ruelle-Bowen measure, which in this case is just the Liouville measure) coincides with the equilibrium state for the null-potential (i.e the measure of maximal entropy). This provides the first direction of Katok’s entropy conjecture, namely: for a locally symmetric metric, the Liouville measure is the measure of maximal entropy.
@@ 2.3.1. Complex hyperbolic quotient
$(M^{n},g)$ being a quotient of the complex hyperbolic space, it is equipped with a complex structure. By this, we mean that there exists a (global) section $J\in C^{\infty}(M;\mathrm{End}(TM))$ satisfying $J^{2}=-\mathrm{Id}$, compatible with the metric structure, i.e $g(\cdot,\cdot)=g(J\cdot,J\cdot)$. Moreover, $(M^{n},g)$ is a Kähler manifold, which means that $J$ is parallel:
@@
see [KN, Chapter IX]. In particular, the parallel vector field $Y_{2}$ appearing in (2.21) can be taken equal to $JY$ in this case.
@@
Since $J$ commutes with the trace, we obtain using (2.6) that $J$ preserves $\Omega_{m}$ for any $m\in\mathbb{N}$. For later use, we also define a twisted version of the symmetrized covariant-derivative:
@@
The proof is the same as in the case of surfaces (see [GuMaz, Proposition 1.51]). The factor $-4$ comes from the fact that for any unit-length vector $v\in T_{x}M$, one has $K_{x}(v,Jv)=-4$, where $K_{x}$ denotes here the sectional curvature, see (2.23).
@@
Here, the raising and lowering operators have the mapping property $\eta^{\delta}_{\varepsilon}:E^{\lambda}_{m}\to E^{\lambda+\delta}_{m+\varepsilon}$ where $\varepsilon,\delta\in\{\pm 1\}$. Finally, if $\lambda>0$, then $\eta_{+}^{+}:E^{\lambda}_{m}\to E^{\lambda+1}_{m+1}$ is injective.
@@
On $\Omega_{0}$, $V$ acts as the zero-operator so its spectrum is equal to $\{0\}$. On $\Omega_{1}$, $V$ acts as $J$ and using $J^{2}=-\mathrm{Id}$, one sees that the spectrum is given by {±$\{\pm i\}$. For $\Omega_{2}$, we compute for any $T\in C^{\infty}(M;S^{2}T^{*}M)$, $V^{2}T=-2T+2T\circ J,$ which implies
@@
In particular, we see that $J$ is a polynomial in $V$ and this implies that the spectrum of $V$ is {±$\{\pm 2i,0\}$. For $\Omega_{3}$, we get this time, for any $T\in C^{\infty}(M;S^{3}T^{*}M)$,
@@
These two equations imply that the spectrum is {±$\{\pm 3i,\pm i\}$.
The mapping properties of $\eta^{\delta}_{\varepsilon}$ is a consequence of the commutation relations (2.32). Let us prove the injectivity statement. For this, we first note that $(\eta_{+}^{+})^{*}=-\eta_{-}^{-}$ and we compute, using $X_{+}^{*}=-X_{-}$ and $H_{+}^{*}=-H_{-}$,
@@
We record the following commutation relations which follows from (2.33):
(2.39) $[\eta_{+}^{+},\eta_{+}^{-}]=0,\quad[\eta_{-}^{+},\eta_{-}^{-}]=0.$
@@
Let $T_{0}\in C^{\infty}(M;S^{m}_{0}T^{*}M)$ and denote by $f=\pi_{m}^{*}T_{0}$ then one has
@@
where in the second bound, we have $m=2$.
@@
2.3.2. Quaternionic hyperbolic quotient.
Quotients of the hyperbolic space $\mathbb{H}\mathbb{H}^{n}$ are examples of a quaternion-Kähler manifolds, [Bes, Chapter 14]. This time, there is no global complex structure but locally, one can find three almost complex structures $J_{1},J_{2},J_{3}:=J_{1}J_{2}$ compatible with the metric $g_{0}$. The endomorphism $J_{i}$ are not defined on the whole manifold but
$Q:=\mathrm{Span}(J_{1},J_{2},J_{3})$
see [BriHa, Proposition 10.12] or [Bes, (14.44)]. Note that it is not clear in this case than one can take $Y_{2}=J_{1}Y,Y_{3}=J_{2}Y$ and $Y_{4}=J_{3}Y$ in (2.21) as the almost complex structure $J_{i}$ may not be parallel but one still has
(2.41) $\mathrm{Span}(Y_{2},Y_{3},Y_{4})=\mathrm{Span}(J_{1}Y,J_{2}Y,J_{3}Y).$
where $U$ is an open set on which $J_{i}$ are defined and span $E$.
2.3.3. Octonionic Cayley plane
This is the last possible type of rank one locally symmetric space of negative curvature. Note that it can only happen in dimension $16$. We refer to the appendix of [Ruan] or [Most, Chapter 19] for more details about the construction. From the remark after [Ruan, Corollary A.6], one sees that there may not exist local endomorphisms $J_{1},\ldots,J_{7}\in\mathrm{End}(TM)$ such that $K_{x}(v,J_{i}v)=-4$ and we will thus use a slightly different approach. Let $\mathbb{O}$ denote the set of octonions. We define the Cayley line through $(a,b)\in\mathbb{O}^{2}\setminus\{0\}$ to be
By [Ruan, Proposition A.5], there exists an isomorphism $\Psi_{x}:\mathbb{O}^{2}\to T_{x}M$ for any point $x\in M$. We define the Cayley line of a vector to be
From [Ruan, Corollary 5.5], we get, with the notations of (2.21), for any $(x,v)\in SM$:

### v3

@@ 2. Preliminaries
In this subsection, we recall some notions and tools from tensor tomography. We refer to [18, Chapter 5.3] for a textbook account of these notions.
@@ 2.1.1. Tensorial analysis
(2.1) $\mathrm{Sym}:V^{\otimes m}\to S^{m}V,\quad e_{i_{1}}\otimes e_{i_{2}}\otimes\cdots\otimes e_{i_{m}}\mapsto\frac{1}{m!}\sum_{\sigma\in\mathfrak{S}_{m}}e_{i_{\sigma(1)}}\otimes e_{i_{\sigma(2)}}\otimes\cdots\otimes e_{i_{\sigma(m)}}$
@@ 2.1.2. Analysis on SMSM
We consider the unit tangent bundle $SM:=\{(x,v)\in TM\mid g(v,v)=1\}$. We will write $\varphi_{t}$ for the geodesic flow. The geodesic vector field is then the generator of the flow, i.e., $X=\tfrac{d}{dt}\varphi_{t}|_{t=0}.$ We denote by $\pi:SM\to M$ the canonical projection on the base. The vertical distribution is defined to be
@@
where $\mathcal{P}_{t}$ denotes the parallel transport along the geodesic. We recall that the metric $g$ on the base induces naturally a metric $g_{\text{Sas}}$ (called the Sasaki metric, see [31, Chapter 1]) on the unit tangent bundle $SM$. We have the following orthogonal splitting:
@@
See [18, Lemma 5.7] for a proof. We see that the conformal factor $\Lambda_{m}^{n}$ depends on the degree $m$ and this will be a central point in the computations of Section 4. In the following, we might suppress the index and write $\langle\cdot,\cdot\rangle$ for the scalar product if it is clear from the context which scalar product is used.
@@
Consider a local frame $(e_{1},e_{2},\ldots,e_{n})$ parallel at $x_{0}$ (i.e., $\nabla_{e_{i}}e_{j}=0$ at $x_{0}$). Recall that the action of $\nabla^{*}$ on a $m$-tensor $T$ can be described locally by:
@@
where $(R(e_{i},v)v)^{\mathbb{V}}\in\mathbb{V}$ is the vertical component of $R(e_{i},v)v$ (in other words, one has $\mathcal{K}(x,v)(R(e_{i},v)v)^{\mathbb{V}}=R(e_{i},v)v$) and the computation of $X,X_{j}]$ can be found in the proof of [27, Lemma 13.1.7]. This means that the expression above simplifies into
@@
We will write $X=X_{+}+X_{-}$, where $X_{\pm}:C^{\infty}(M,\Omega_{k})\to C^{\infty}(M,\Omega_{k\pm 1})$, see [27, Lemma 14.2.1].
@@
We have the important relation, see [27, Lemma 14.1.8]:
@@ 2.1.4. Pestov’s identity
We introduce an “energy estimate” due to Pestov (and first written in coordinate-free way in [25, Appendix Theorem 1.1], see also [32, Corollary 3.6]). It will be crucial in the proof of solenoidal injectivity of Proposition 4.1. More precisely, we will need the following estimate which is implied in negative curvature by the localized Pestov identity. We will use the result of [27, Theorem 14.3.4, Equation (16.1.3)], which is valid in dimension $n\geq 4$:
@@
A consequence is that the operator $X_{+}$ is injective on $\Omega_{m}$ for $m\geq 1$ ([27, Lemma 16.1.6]):
@@
Finally, let $S\in C^{\infty}(M;S^{2}T^{*}M)$ be a divergence-free tensor, i.e., $D_{g_{0}}^{*}S=0$ and write $S=S_{0}+L(h)$ with $S_{0}\in C^{\infty}(M;S_{0}^{2}T^{*}M)$ and $h\in C^{\infty}(M)$. Then
$D_{g_{0}}^{*}S=0\iff D_{g_{0}}^{*}(S_{0}+L(h))=0\iff D_{g_{0}}^{*}S_{0}-D_{g_{0}}h=0,$
where we have used the relation $D_{g_{0}}^{*}L(h)=D_{g_{0}}h$ on 0-tensors. The divergence operator corresponds to a rescaled version of $X_{-}$. Indeed, using (2.10) and (2.7), one has, for any $S_{0}\in C^{\infty}(M;S^{2}_{0}T^{*}M)$ and $p\in C^{\infty}(M;T^{*}M)$,
@@
In particular, the functional $\Phi$ is constant along any orbit $\mathcal{O}(g)$ and the kernel of the Hessian $d^{2}\Phi_{g_{0}}$ always contains the tangent space $T_{g_{0}}\mathcal{O}(g_{0})$. We will hence need to prove an injectivity result of the Hessian on a “transverse slice” to $T_{g_{0}}\mathcal{O}(g_{0})$. We remark the following fact:
@@
in particular, a natural transverse slice is provided by the kernel of the adjoint $D_{g_{0}}^{*}$, [27, Theorem 14.1.10].
@@
The analysis will take place on the space of divergence-free (or solenoidal) tensors $C^{\infty}(M;S^{m}T^{*}M)\cap\mathrm{Ker}(D_{g_{0}}^{*})$. We thus recall the following lemma which allows to “project” a metric $g$ on solenoidal tensors. It was obtained in this form in [16, Lemma 2.4] but the idea goes back to Ebin [9].
@@
Let $k\geq 2$, and $\alpha\in(0,1)$. Then there exists a neighborhood $\mathcal{U}$ of $g_{0}$ in the $C^{k,\alpha}$-topology such that for any $g\in\mathcal{U}$, there is a unique $\phi\in\mathrm{Diff}_{0}(M)$ of regularity $C^{k+1,\alpha}$, close to identity, such that $\phi^{*}g\in\mathrm{Ker}(D_{g_{0}}^{*})$ is divergence free. Moreover, there exists $\epsilon>0$ and $C>0$ such that
$\|g-g_{0}\|_{C^{k,\alpha}}\leq\epsilon\ \Rightarrow\ \|\phi^{*}g-g_{0}\|_{C^{k,\alpha}}\leq C\|g-g_{0}\|_{C^{k,\alpha}}.$
@@ 2.2. The operator $\Pi$
We recall in this subsection the main feature of the “generalized X-ray transform” $\Pi$. It was first introduced by Guillarmou in [15] and used by Guillarmou and Lefeuvre in [17] for their proof of the local injectivity of the marked length spectrum. A quick definition of the operator can be given using the variance. Let $f$ have zero mean, i.e., $\int_{SM}f(z)d\mu_{\mathrm{Liou}}^{g}(z)=0$. The variance of $f$, with respect to the Liouville measure $\mu^{g}_{\mathrm{Liou}}$ is
(2.15) $\mathrm{Var}_{\mu^{g}_{\mathrm{Liou}}}(f):=\lim_{T\to+\infty}\frac{1}{T}\int_{SM}\left(\int_{0}^{T}f(\varphi_{t}(z))dt\right)^{2}d\mu_{\mathrm{Liou}}^{g}(z).$
Then for any smooth function $f\in C^{\infty}(SM)$, we define
(2.16) $\quad\langle\Pi f,f\rangle=\int_{\mathbb{R}}\langle f\circ\varphi_{t},f\rangle_{L^{2}}dt=\mathrm{Var}_{\mu_{\mathrm{Liou}}^{g}}(f).$
The second equality follows from the exponential decay of correlations [28], see for instance [16, Equation $(2.6)$]. The Hessian $d^{2}\Phi(g_{0})$ will be rewritten using the operator $\Pi$ in Theorem 5.
A more microlocal definition of the operator $\Pi$ was given originally in [15] which allowed to track its analytical properties but we will not need to recall this definition here. We now define
@@
The operator $\Pi_{m}$ was shown to be a pseudo-differential operator of order 1-1 which is elliptic and injective on solenoidal tensors (see [15, 17, Theorem 3.5, Lemma 4.3]). For any $Q\in C^{\infty}(M;S^{m}T^{*}M)$, using (2.16) gives
(2.18) $\mathrm{Var}_{\mu_{\mathrm{Liou}}^{g_{0}}}(\pi_{m}^{*}Q,\pi_{m}^{*}Q)=\langle\Pi_{2}^{g_{0}}Q,Q\rangle-\langle\pi_{2}^{*}Q,1\rangle_{L^{2}(SM)}^{2}.$
In negative sectional curvature, we have the following coercive estimate, which we will use crucially to obtain the stability estimate (1.3) [16, Lemma 2.1].
@@ 2.3. Locally symmetric spaces
Let $M^{n}$ be a differentiable manifold. A metric $g$ on $M$ is locally symmetric if its curvature tensor $R_{g}$ is parallel : $\nabla_{g}R_{g}=0$.
If $(x,v)\in SM$ and $Y$ is the corresponding geodesic vector field, then one can find parallel orthonormal fields $Y_{1}=Y,Y_{2},\ldots,Y_{n}$ (along the geodesic defined by $(x,v)$) such that $R(Y,Y_{j})Y=-\lambda_{j}^{2}Y_{j}$ for some $\lambda_{j}$ that do not depend on $(x,v)$.
This means that the unstable Jacobian $J^{u}(v):=-\tfrac{d}{dt}|_{t=0}\mathrm{det}(d\varphi_{t}(v)|_{E_{u}(v)})$, where $E_{u}(v)$ is the unstable bundle at $v\in SM$, is a constant function of $v$. In particular, its equilibrium state (the Sinai-Ruelle-Bowen measure, which in this case is just the Liouville measure) coincides with the equilibrium state for the null-potential (i.e., the measure of maximal entropy). This provides the first direction of Katok’s entropy conjecture, namely: for a locally symmetric metric, the Liouville measure is the measure of maximal entropy.
There is a more algebraic characterization of locally symmetric spaces. Indeed, it is known that [30, Chapter 19] compact locally symmetric spaces of negative sectional curvature are obtained as compact quotients of hyperbolic spaces. More precisely, one has
@@
We note that for a locally symmetric metric, the Lyapunov exponents of the geodesic flow are given by:
@@
In particular, we have
@@
where $Z\mapsto R(Y_{j},Z)Y_{j}$ is a symmetric endomorphism. Moreover, since one has $4\leq g(R(Y_{j},Z)Y_{j},Z)\leq-1$ for any unit length $Z$, this implies that
@@ 2.3.1. Complex hyperbolic quotient
Since $(M^{n},g)$ is a quotient of the complex hyperbolic space, it is equipped with a complex structure. By this, we mean that there exists a (global) section $J\in C^{\infty}(M;\mathrm{End}(TM))$ satisfying $J^{2}=-\mathrm{Id}$, compatible with the metric structure, i.e., $g(\cdot,\cdot)=g(J\cdot,J\cdot)$. Moreover, $(M^{n},g)$ is a Kähler manifold, which means that $J$ is parallel:
@@
see [34, Chapter IX]. In particular, the parallel vector field $Y_{2}$ appearing in (2.22) can be taken equal to $JY$ in this case.
@@
Since $J$ commutes with the trace, we obtain using (2.6) that $J$ preserves $\Omega_{m}$ for any $m\in\mathbb{N}$. For later use, we also define a twisted version of the symmetrized covariant derivative:
@@
2.3.2. Quaternionic hyperbolic quotient.
Quotients of the hyperbolic space $\mathbb{H}\mathbb{H}^{n}$ are examples of a quaternion-Kähler manifolds, [26, Chapter 14]. This time, there is no global complex structure but locally, one can find three almost complex structures $J_{1},J_{2},J_{3}:=J_{1}J_{2}$ compatible with the metric $g_{0}$. The endomorphism $J_{i}$ are not defined on the whole manifold but
$Q:=\mathrm{Span}(J_{1},J_{2},J_{3})$
see [3, Proposition 10.12] or [26, (14.44)]. Note that it is not clear in this case than one can take $Y_{2}=J_{1}Y,Y_{3}=J_{2}Y$ and $Y_{4}=J_{3}Y$ in (2.22) as the almost complex structure $J_{i}$ may not be parallel but one still has
(2.28) $\mathrm{Span}(Y_{2},Y_{3},Y_{4})=\mathrm{Span}(J_{1}Y,J_{2}Y,J_{3}Y).$
where $U$ is an open set on which $J_{i}$ are defined and span $E$.
2.3.3. Octonionic Cayley plane
This is the last possible type of rank one locally symmetric space of negative curvature. Note that it can only happen in dimension $16$. We refer to the appendix of [33] or [30, Chapter 19] for more details about the construction. From the remark after [33, Corollary A.6], one sees that there may not exist local endomorphisms $J_{1},\ldots,J_{7}\in\mathrm{End}(TM)$ such that $K_{x}(v,J_{i}v)=-4$ and we will thus use a slightly different approach. Let $\mathbb{O}$ denote the set of octonions. We define the Cayley line through $(a,b)\in\mathbb{O}^{2}\setminus\{0\}$ to be
By [33, Proposition A.5], there exists an isomorphism $\Psi_{x}:\mathbb{O}^{2}\to T_{x}M$ for any point $x\in M$. We define the Cayley line of a vector to be
From [33, Corollary 5.5], we get, with the notations of (2.22), for any $(x,v)\in SM$:
2.4. A theory of raising and lowering operators for complex hyperbolic spaces.
In this subsection, we build an analogous theory of the Guillemin–Kazhdan [19] raising/lowering operators for surfaces for complex hyperbolic spaces. For a textbook account of the theory for surfaces, see [27, Chapter 15.2] or [18, Chapter 1.11]. The key point is that the complex structure recalled in (2.23) allows for a further splitting of the space of spherical harmonics. In particular, the geodesic vector field $X$ splits into raising and lowering operators which have specific mapping properties with respect to this splitting. In Lemma 2.7, we show that these operators are injective when restricted to certain spaces and this will be the key of the proof of Theorem 3 in Section 4.1.
@@ 2.3.3. Octonionic Cayley plane
The proof is the same as in the case of surfaces (see [18, Proposition 1.51]). The factor 4-4 comes from the fact that for any unit-length vector $v\in T_{x}M$, one has $K_{x}(v,Jv)=-4$, where $K_{x}$ denotes here the sectional curvature, see (2.24).
@@
In the cases of surfaces, these operators were introduced and studied by Guillemin and Kazhdan, see [19, Section 3].
@@
Here, the raising and lowering operators have the mapping property $\eta^{\delta}_{\epsilon}:E^{\lambda}_{m}\to E^{\lambda+\delta}_{m+\epsilon}$ where $\epsilon,\delta\in\{\pm 1\}$. Finally, if $\lambda>0$, then $\eta_{+}^{+}:E^{\lambda}_{m}\to E^{\lambda+1}_{m+1}$ is injective.
@@
On $\Omega_{0}$, $V$ acts as the zero-operator so its spectrum is equal to $\{0\}$. On $\Omega_{1}$, $V$ acts as $J$ and using $J^{2}=-\mathrm{Id}$, one sees that the spectrum is given by $\{\pm i\}$. For $\Omega_{2}$, we compute for any $T\in C^{\infty}(M;S^{2}T^{*}M)$, $V^{2}T=-2T+2T\circ J,$ which implies
@@
In particular, we see that $J$ is a polynomial in $V$ and this implies that the spectrum of $V$ is $\{\pm 2i,0\}$. For $\Omega_{3}$, we get this time, for any $T\in C^{\infty}(M;S^{3}T^{*}M)$,
@@
These two equations imply that the spectrum is $\{\pm 3i,\pm i\}$.
The mapping properties of $\eta^{\delta}_{\epsilon}$ is a consequence of the commutation relations (2.39). Let us prove the injectivity statement. For this, we first note that $(\eta_{+}^{+})^{*}=-\eta_{-}^{-}$ and we compute, using $X_{+}^{*}=-X_{-}$ and $H_{+}^{*}=-H_{-}$,
@@
We record the following commutation relations which follow from (2.40):
(2.46) $\eta_{+}^{+},\eta_{+}^{-}]=0,\quad[\eta_{-}^{+},\eta_{-}^{-}]=0.$
@@
Let $T_{0}\in C^{\infty}(M;S^{m}_{0}T^{*}M)$ and denote by $f=\pi_{m}^{*}T_{0}$. Then one has
@@
where in the second bound, we have $m=2$ and $\Lambda_{m}^{n}$ is defined in 2.7.